OpenStudy (anonymous):

What is the approximate distance from (-2,3) to (4,1)?

OpenStudy (saifoo.khan):

6.32456

OpenStudy (anonymous):

Based on the Pythagorean theorem: \[d = \sqrt{(x_2-x_1)^2 + (y_2 - y_1)^2}\] The original statement: \[x^2 + y^2 = d^2\]

OpenStudy (anonymous):

6.234

OpenStudy (anonymous):

\[\sqrt{(4 - (-2))^2 + (1 - 3)^2} = \sqrt{(6)^2 + (-2)^2} = \sqrt{40} = 2\sqrt{10} \approx 6.3245553203\]